EVA Report – November 27th

Crew 216 EVA Report 27-Nov-19
EVA # 05
Author: Levesque
Purpose of EVA: To explore suitability for possible radio repeater sites close to the Hab. Collect footage for documentary.
Start time for EVA: 0930
End time for EVA: 1130
Narrative: EVA #5 was our crew’s first non-rover EVA. We left the Airlock at 0930 with Levesque, Lawson, Alexandrova, and Kennedy. We traveled west up the bentonite hills adjacent to the Hab, stopping a few times for footage and oxygen absorption. At the summit of the ridge we were greeted by the majestic views in all directions, looking down onto our decorous Hab below. We then traveled south to the coordinate #1. We observed a surprising amount of fossils for Mars, but found the road in good condition and excellent for travel. From location #1 we stopped, did some filming, and returned north on the road we had been traveling. We arrived at the coordinate #2 shortly after. This is the site of the old repeater according to our documents. We evaluated the suitability of the location. It would do well for comms east west and south, but site #3 would most likely provide a better overall coverage, especially north. We did some more filming with only minor queues from our documentarian and continued north on the Hab Ridge road. Shortly after we came to the break in the ridge and the old access road. We followed this down until we were at the foot of the north ridge, the base of the hill of coordinate #3. We were concerned that with the time we had left that it would be cutting it close, so we elected to reevaluate the location at a later EVA and return to the Hab. We did some minor filming and savored how we are living our best lives. We did a quick external evaluation of the Hab and returned to the airlock approximately at 1130.
Destination: In the vicinity of Hab (within 1 km), 3 specific points listed below.
Coordinates (use UTM NAD27 CONUS):
1: 12 S 0517874 4250342.
2: 12 S 0517887 4250342.
3: 12 S 0518464 4251538.
EVA Participants: Levesque, Lawson, Alexandrova, Kennedy
Road(s) and routes per MDRS Map:
1) Hab to 12 S 0517874 4250342 on foot
2) 12 S 0517874 4250342 to 12 S 0517887 4250342 on foot
3) 12 S 0517887 4250342 to 12 S 0518464 4251538 on foot
4) 12 S 0518464 4251538 to Hab on foot
5) Total distance 3.5 km
Mode of travel: Walking with powerful self determination
Vehicles used: N/A

EVA Report – November 27th

Crew 216 EVA Report 27-Nov-19
EVA # 05
Author: Levesque
Purpose of EVA: To explore suitability for possible radio repeater sites close to the Hab. Collect footage for documentary.
Start time for EVA: 0930
End time for EVA: 1130
Narrative: EVA #5 was our crew’s first non-rover EVA. We left the Airlock at 0930 with Levesque, Lawson, Alexandrova, and Kennedy. We traveled west up the bentonite hills adjacent to the Hab, stopping a few times for footage and oxygen absorption. At the summit of the ridge we were greeted by the majestic views in all directions, looking down onto our decorous Hab below. We then traveled south to the coordinate #1. We observed a surprising amount of fossils for Mars, but found the road in good condition and excellent for travel. From location #1 we stopped, did some filming, and returned north on the road we had been traveling. We arrived at the coordinate #2 shortly after. This is the site of the old repeater according to our documents. We evaluated the suitability of the location. It would do well for comms east west and south, but site #3 would most likely provide a better overall coverage, especially north. We did some more filming with only minor queues from our documentarian and continued north on the Hab Ridge road. Shortly after we came to the break in the ridge and the old access road. We followed this down until we were at the foot of the north ridge, the base of the hill of coordinate #3. We were concerned that with the time we had left that it would be cutting it close, so we elected to reevaluate the location at a later EVA and return to the Hab. We did some minor filming and savored how we are living our best lives. We did a quick external evaluation of the Hab and returned to the airlock approximately at 1130.
Destination: In the vicinity of Hab (within 1 km), 3 specific points listed below.
Coordinates (use UTM NAD27 CONUS):
1: 12 S 0517874 4250342.
2: 12 S 0517887 4250342.
3: 12 S 0518464 4251538.
EVA Participants: Levesque, Lawson, Alexandrova, Kennedy
Road(s) and routes per MDRS Map:
1) Hab to 12 S 0517874 4250342 on foot
2) 12 S 0517874 4250342 to 12 S 0517887 4250342 on foot
3) 12 S 0517887 4250342 to 12 S 0518464 4251538 on foot
4) 12 S 0518464 4251538 to Hab on foot
5) Total distance 3.5 km
Mode of travel: Walking with powerful self determination
Vehicles used: N/A

Research Report – November 27th

Crew 216 Research Report 27-Nov-2019
Sol 3
Submitted by: Crew Researcher Whittle

1) Psych Scale (TAMU IRB2019-1042D) – Next surveys planned for Sol 5.

2) Medical Casts – Continuation of Crew 215 research. Planning to crush gypsum tomorrow.

3) In-Situ Fluorescent Mineral Prospecting – Still searching for radioactives and rare Earths, looking at potential EVA plans for week 2.

4) MDRS Mapping – Nothing further to report (awaiting return of map with desired annotations from Director).

5) MDRS Communications – Completed 2 EVAs to evaluate potential rebroadcasting sites. GPS coordinates of 5 potential sites marked (3 close to Hab, 2 on Kissing Camel Ridge). Plan to evaluate further with GIS.

Operations Report – November 27th

Crew 216 Operations Report 27-Nov-19
SOL: 3
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue.
Notes on non-nominal systems: Outpost directed Opportunity to be charged at Spirit’s
traditional port.
Generator: run
Hours run: 12.1
From what time last night: 1850
To what time this morning: 0658
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 55%
Diesel Reading – 41
Station Propane Reading – 59
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1461926
Water (static tank) (gal): 401
Static to Loft Pump used – yes
Water in Green Hab (gal): 272
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: yes
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 81
Currently charging: no (see notes non-nominal systems).
Opportunity rover used: no
Hours: 68.7
Beginning charge: 46
Ending charge: 46
Currently charging: yes (in Spirit port)
Curiosity rover used: yes
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 81
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments:
Summary of internet: No serious issues.
Summary of suits and radios: Midland GXT serial # A1707006944, erratic power setting fluctuations,
possibly intermittent battery connectivity while on afternoon EVA.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: UV scan of gypsum samples performed. In addition to gypsum, samples
proabably include calcite and selenite.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues.
Summary of any observatory issues: Visual observation of sun AM.
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: Two bags

Operations Report – November 27th

Sorry about the Report! This doesn’t belong to this e-list.

On Fri, Nov 29, 2019 at 7:44 PM Atila Meszaros <atila> wrote:

Crew 216 Operations Report 27-Nov-19
SOL: 3
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue.
Notes on non-nominal systems: Outpost directed Opportunity to be charged at Spirit’s
traditional port.
Generator: run
Hours run: 12.1
From what time last night: 1850
To what time this morning: 0658
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 55%
Diesel Reading – 41
Station Propane Reading – 59
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1461926
Water (static tank) (gal): 401
Static to Loft Pump used – yes
Water in Green Hab (gal): 272
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: yes
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 81
Currently charging: no (see notes non-nominal systems).
Opportunity rover used: no
Hours: 68.7
Beginning charge: 46
Ending charge: 46
Currently charging: yes (in Spirit port)
Curiosity rover used: yes
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 81
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments:
Summary of internet: No serious issues.
Summary of suits and radios: Midland GXT serial # A1707006944, erratic power setting fluctuations,
possibly intermittent battery connectivity while on afternoon EVA.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: UV scan of gypsum samples performed. In addition to gypsum, samples
proabably include calcite and selenite.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues.
Summary of any observatory issues: Visual observation of sun AM.
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: Two bags

Operations Report – November 27th

Crew 216 Operations Report 27-Nov-19
SOL: 3
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue.
Notes on non-nominal systems: Outpost directed Opportunity to be charged at Spirit’s
traditional port.
Generator: run
Hours run: 12.1
From what time last night: 1850
To what time this morning: 0658
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 55%
Diesel Reading – 41
Station Propane Reading – 59
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1461926
Water (static tank) (gal): 401
Static to Loft Pump used – yes
Water in Green Hab (gal): 272
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: yes
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 81
Currently charging: no (see notes non-nominal systems).
Opportunity rover used: no
Hours: 68.7
Beginning charge: 46
Ending charge: 46
Currently charging: yes (in Spirit port)
Curiosity rover used: yes
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 81
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments:
Summary of internet: No serious issues.
Summary of suits and radios: Midland GXT serial # A1707006944, erratic power setting fluctuations,
possibly intermittent battery connectivity while on afternoon EVA.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: UV scan of gypsum samples performed. In addition to gypsum, samples
proabably include calcite and selenite.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues.
Summary of any observatory issues: Visual observation of sun AM.
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: Two bags

GreenHab Report – November 27th

Crew 216 GreenHab Report 27-NOV-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 12.15pm, closed at 4.40 pm.
Average temperature: 17C
Hours of supplemental light: Automatic system
Daily water usage for crops: 13L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 10.00 am; 4.30 pm.
Sprouted:
from both long trays of narcissis,
from both trays of cherry tomatoes,
from all 3 round trays of snow peas.

SEASONAL PLANTING HISTORY:
(Established plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)

OTHER NOTES
Harvest: None
Support/supplies requested: None

Journalist Report – November 27th

Crew 216 Journalist Report 27-NOV-19
Sol 3
Author: Evgenia Alexandrova

Light goes up and the water pump notifies you that the new day on Mars has just began. You have your dehydrated breakfast and hurry to spend next 20 minutes dressing up and preparing the gear for an EVA. Here you are: spacesuit on with the ventilation whispering in your ear and your back telling you why on Earth (oops on Mars) would you want to go out today. You spend 40 minutes climbing with your spacesuit that hides your feet from your sight of view, with your gear (and here you start considering downshifting) and you slide, and you almost fall. Once on top you can’t even breath, neither see anything because of the fog from your breath. And suddenly… You get a stroke. A happiness stroke. You see that view and just realize you are exactly where you wished you were at that very moment. Happiness is a choice. And ARES 216 chose MDRS.

So our EVA was quite enough

To the top of the ridge it was tough

She filmed us all three

For a documentree

And she carried a boat load of stuff

Journalist Report – November 22th

Crew 216 Journalist Report 27-NOV-19
Sol 3
Author: Evgenia Alexandrova

Light goes up and the water pump notifies you that the new day on Mars has just began. You have your dehydrated breakfast and hurry to spend next 20 minutes dressing up and preparing the gear for an EVA. Here you are: spacesuit on with the ventilation whispering in your ear and your back telling you why on Earth (oops on Mars) would you want to go out today. You spend 40 minutes climbing with your spacesuit that hides your feet from your sight of view, with your gear (and here you start considering downshifting) and you slide, and you almost fall. Once on top you can’t even breath, neither see anything because of the fog from your breath. And suddenly… You get a stroke. A happiness stroke. You see that view and just realize you are exactly where you wished you were at that very moment. Happiness is a choice. And ARES 216 chose MDRS.

So our EVA was quite enough

To the top of the ridge it was tough

She filmed us all three

For a documentree

And she carried a boat load of stuff